5y+4y=19
y=6x-3
solve by the substitution method and by showing the steps.

Answer by mananth(16949)

5y+4y=19
y=6x-3
9y=19
substitute y in 9y=19
9(6x-3)=19
54x-27=19
+27
54x=46
x=46/54
x=23/27
..
m.ananth@hotmail.ca
..
y=(6*23/27)-3
y=138/27 -3
y=(138-81)/27
y=57/27
y=19/9
...
m.ananth@hotmail.ca